OpenStudy (anonymous):

A 1,700 kg car coasts down a hill starting from rest and decreases its height by 22 m over a distance of 110 m of travel. How fast will it be going at the end of the hill, if friction has no effect in slowing its motion?

OpenStudy (anonymous):

Hi Welcome to Openstudy! I think you can use Energy conservation here: let's see... Ek= kinetic energy = 1/2mv^2 Ep= potential energy = mgh given: m= 1700 kg vi=0m/s di = 110 m //decreases by 22 m vf= ? when df = 0 accdg to the law: Total initial mechanical energy = total Final mechanical energy \(\sf E_{ki} + E_{pi} = E_{kf} + E_{pi}\\1/2mv_i^2+mgd_i=1/2mv_f^2+mgd_f\\0+mgd_i=1/2mv_f^2+0\\v_f= \sqrt{2gd_i}\)